OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). 2 sin2^x = sin x

OpenStudy (anonymous):

\[2\sin 2^{x}=\sin x \] are you sure it looks like that?

OpenStudy (anonymous):

\[2 \sin ^{2}x=\sin x\]

OpenStudy (anonymous):

sorry this is how it looks like

OpenStudy (anonymous):

pi / 4 = 45 5pi / 4 = 135

OpenStudy (anonymous):

2 sin^2 x = sin x 2 sin x = 1 dividing both sides by sin x sin x = 1 / 2 dividing both sides by 2 arcsin (1/2) = x so x = pi/4, 5pi/4